2 passenger trains traveling in opposite directions meet and pass each other. Each train is 1/12 mi long and is traveling at 50 mph. How many seconds after front part of the trains meet will their rear parts pass each other?
Answer by CharlesG2(834)

2 passenger trains traveling in opposite directions meet and pass each other. Each train is 1/12 mi long and is traveling at 50 mph. How many seconds after front part of the trains meet will their rear parts pass each other?

1/12 mile * 5280 feet/mile = 440 feet per train
50 miles/hour * 1/3600 hr/sec * 5280 ft/mile = 73 1/3 ft/second
one train has to travel 440 feet to pass the other train
440 feet * 1/(73 1/3) sec/ft = 6 seconds